Equations Inequalities and applications
(2/3)x + 4 = (3/4)x + 2
???

Answer by lordacro(12)

(2/3)x + 4 = (3/4)x + 2
First what you want to do is put like things on one side. For example, put all the X's on one side and all numbers on the other.
Let's subtract (2/3)x from both sides. This means you have to do 3/4 minus 2/3 to find out how many X's you have on the right. The answer to that is 1/12.
Now you have
4 = (1/12)x + 2
Now put your numbers on the other side. You want to get rid of the 2, so subtract it from both sides.
Now you have
2 = (1/12)x
Or "two is one-twelfth of some number." If you multiply both sides by 12, you will find what x is.
2(12) = x
24 = x